OpenStudy (anonymous):

Simplify 2 square root of 108

OpenStudy (mertsj):

\[2\sqrt{108}=2\sqrt{36(3)}=\]

OpenStudy (mertsj):

Can you finish it now?

OpenStudy (anonymous):

is it 6 square root of 3?

OpenStudy (mertsj):

Did you forget about the 2 out in front?

OpenStudy (anonymous):

oh! 12 square root of 3

OpenStudy (mertsj):

\[2\sqrt{36(3)}=2\sqrt{36}\sqrt{3}=2(6)\sqrt{3}=12\sqrt{3}\]